A garden supply store sells Two types of lawn mowers. Total sales of mowers for the year were $8379.70. The total number of mowers sold was 30. The small mower costs $249.99. The large mower costs $329.99. Find the number sold of each type of mower.
Answer by mananth(16949)

small nower x numbers
large mower y numbers
...
x+y=30 .............1
249.99x+ 329.99y =8379.7 .............2

multiply (1)by -329.99
Multiply (2) by 1
-329.99 x-329.99 y=-9899.7
249.99x +329.99y =8379.7

Add the two equations
-80x=-1520
/-80
x=19 small mowers
plug value of x in (1)
x+y=30
19+y=30
y=30-19
y=11 larger mowers